A particle of mass 3m is projected from the ground at an angle with the horizontal the range is R. At the highes point o itsvpah it breaks into two pieces  m and  2m. tthe smaller mass stops  and larger mass falls at a distance x from point of projection where x=?

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
3m breaks into 2 parts at highest point (R/2) 
Now the smaller one (m) falls at R/2 and 2m keep flight and it goes x from point of projection but the center of mass should lend at R
So m(R/2)+2m(x)=3m(R) => 2m(x)=(5/2)mR=> x=5/4R Ans